Derive an expression for adiabatic lapse
rate.
1
Expert's answer
2019-10-28T10:55:07-0400

We will now derive an expression for the rate of change of temperature with height of a parcel of dry air as it moves about in the Earth’s atmosphere. Since the air parcel undergoes only adiabatic transformations


"dQ=0"


and the atmosphere is in hydrostatic equilibrium.

The gravity side


"dp=-g\cdot\rho\cdot dz"


"\frac{dp}{\rho}=-g\cdot dz"


The thermodynamics side


"-dA=dU \to-pdV=mc_{V}dT"


"Vdp-d(pV)=mc_{V}dT"


"Vdp-d(nRT)=Vdp-nRdT=mc_{V}dT"


"\frac{V}{m}dp-\frac{n}{m}RdT=c_{V}dT"


"\frac{1}{\rho}dp-\frac{n}{nM}RdT=c_{V}dT"


"\frac{1}{\rho}dp=(\frac{R}{M}+c_{V})dT"


"\frac{dp}{\rho}=c_{p}dT"


"c_{p}dT=-g\cdot dz"


And finally


"\frac{dT}{dz}=-\frac{g}{c_{p}}".